### v1

# A quantitative Robbins-Siegmund theorem

Abstract.
The Robbins-Siegmund theorem is one of the most important results in stochastic optimization, where it is widely used to prove the convergence of stochastic algorithms. We provide a quantitative version of the theorem, establishing a bound on how far one needs to look in order to locate a region of metastability in the sense of Tao. Our proof involves a metastable analogue of Doob’s theorem for $L_{1}$-supermartingales along with a series of technical lemmas that make precise how quantitative information propagates through sums and products of stochastic processes. In this way, our paper establishes a general methodology for finding metastable bounds for stochastic processes that can be reduced to supermartingales, and therefore for obtaining quantitative convergence information across a broad class of stochastic algorithms whose convergence proof relies on some variation of the Robbins-Siegmund theorem. We conclude by discussing how our general quantitative result might be used in practice.
1. Introduction
1.1. The Robbins-Siegmund theorem and its significance
The monotone convergence theorem, by which we mean the statement that any monotone, bounded sequence of real numbers converges, is a basic fact of real analysis. Yet variants of the monotone convergence theorem, specifically results pertaining to convergence properties of almost-monotone sequences, are central to optimization, fixed-point theory, game theory, and related areas. Consider the following example:
Theorem 1.1 (Lemma 5.31 of [4]).
Let $\{{x_{n}}\}$, $\{{\alpha_{n}}\}$, $\{{\beta_{n}}\}$ and $\{{\gamma_{n}}\}$ be sequences of nonnegative reals with $\sum_{i=0}^{\infty}\alpha_{i}<\infty$ and $\sum_{i=0}^{\infty}\gamma_{i}<\infty$ such that
$x_{n+1}\leq(1+\alpha_{n})x_{n}-\beta_{n}+\gamma_{n}$
for all $n\in\mathbb{N}$. Then $\{{x_{n}}\}$ converges and $\sum_{i=0}^{\infty}\beta_{i}<\infty$.
Though completely elementary in appearance, Theorem 1.1 represents a crucial lemma for numerous convergence proofs in fixed-point theory: See [48] for the special case $\beta_{n}=0$, and [4, 13] for further examples and references. In concrete applications, we might have $x_{n}:=\lVert{v_{n}-p}\rVert$ for an iterative algorithm $\{{v_{n}}\}$ in a normed space, $p$ a solution to some problem at hand (such as a fixed point of an operator $T$), $\{{\alpha_{n}}\}$ the step-sizes of an algorithm and $\{{\gamma_{n}}\}$ some error terms. The convergence of $\lVert{v_{n}-p}\rVert$ is then in turn usually a stepping stone towards proving something more concrete that justifies the algorithm, such as $\lVert{v_{n}-p}\rVert\to 0$ or $\lVert{Tv_{n}-v_{n}}\rVert\to 0$.
The scenario sketched above is just one of a myriad of examples where abstract convergence results on almost-monotone sequences of real numbers have found applications in different areas of functional analysis and optimization. An extensive survey of this general proof strategy can be found in [13], where Table 1 of that paper gives a good impression of just how widely applicable this strategy has been (and Theorem 1.1 – which appears there as Lemma 3.6 – is one in a list of 22 such abstract convergence results for deterministic sequences).
Interestingly but unsurprisingly, an entirely analogous phenomenon appears in the field of stochastic approximation, where roughly speaking, monotone convergence is replaced by martingale convergence, and now abstract convergence results relating to almost-supermartingales are of fundamental importance. Among these, the most well-known is the so-called Robbins-Siegmund theorem, a direct stochastic analogue of Theorem 1.1:
Theorem 1.2 (Robbins-Siegmund [50]).
Let $\{{X_{n}}\}$, $\{{A_{n}}\}$, $\{{B_{n}}\}$ and $\{{C_{n}}\}$ be sequences of nonnegative integrable random variables on some arbitrary probability space and adapted to the filtration $\{{\mathcal{F}_{n}}\}$, with $\sum_{i=0}^{\infty}A_{i}<\infty$ and $\sum_{i=0}^{\infty}C_{i}<\infty$ almost surely and
$\mathbb{E}[X_{n+1}\mid\mathcal{F}_{n}]\leq(1+A_{n})X_{n}-B_{n}+C_{n}$
almost surely for all $n\in\mathbb{N}$. Then, almost surely, $\{{X_{n}}\}$ converges and $\sum_{i=0}^{\infty}B_{i}<\infty$.
The Robbins-Siegmund theorem has been used to establish the convergence of a broad range of classic stochastic approximation algorithms, most famously those based on the Robbins-Monro scheme [49], such as stochastic gradient descent. In this case in particular, the two parts of the theorem are used in tandem to prove the stronger result $X_{n}\to 0$ subject to additional conditions on the parameters, where now we might have $X_{n}:=\lVert{v_{n}-\theta}\rVert^{2}$ for $\{{v_{n}}\}$ the stochastic algorithm and $\theta$ the root of some regression function (see [50] or the survey [39] for examples of this).
Given the significance of gradient descent methods in modern artificial intelligence, the Robbins-Siegmund theorem continues to be widely used, though its importance is by no means restricted to methods based on Robbins-Monro: Theorem 1.2 is generally applicable to any kind of stochastic process that enjoys a quasi-Fejér monotonicity property, of which there are many examples (cf. the early paper [12] and the more recent developments along these lines in [6]).
1.2. Basic idea of the paper
When simple abstract convergence results like Theorems 1.1 and 1.2 have so many concrete applications, it is natural to ask whether we can strengthen those abstract results with additional information, in the hope that this information is also relevant for the applications. One of the most obvious forms such additional information might take is numerical information on the speed of convergence. It has been repeatedly shown that tools from mathematical logic are adept at both obtaining this sort of quantitative information and exploiting it in concrete scenarios, to the extent that an entire subfield of logic is now devoted to this aim, the so-called proof mining program. This paper constitutes one of the first such applications of logic in stochastic optimization, where our focus in on the Robbins-Siegmund theorem and related results.
1.3. Proof mining and stochastic algorithms
Proof mining, or applied proof theory, refers to a large body of work – originating with the Kreisel [36, 37] in the 50s and brought to maturity from the 90s by Kohlenbach and his collaborators [25] – that uses insights and techniques from logic to obtain both quantitative and qualitative improvements of theorems in “mainstream” mathematics through the analysis of their proofs. A typical result here is the extraction of computational information from a seemingly nonconstructive proof that is both of low (typically polynomial) complexity, and highly uniform in that it is independent of the precise structure of the underlying spaces. In the process, one is often able to generalise the proof, for example to a broader class of spaces, or by demonstrating that an assumption of the original theorem can be weakened. Alongside these applications lie so-called logical metatheorems (starting with [24] and the most recent in [42]), which explain those applications as instances of general logical phenomena and help guide the search for future breakthroughs.
Over the last decades, proof mining has been hugely successful in the nonstochastic setting, particularly in nonlinear analysis, fixed-point theory, convex optimization and related areas (see the survey papers [26, 27] for an overview of more recent work). In many cases, there is a similar reliance on quantitative convergence lemmas for almost-monotone sequences, and typically these quantitative lemmas then find further applications in future work (see [43] for a recent survey on the analysis of abstract convergence results from the quantitative perspective). For example, a quantitative variant of a special case of Theorem 1.1 was first given in [29] in order to study asymptotically nonexpansive mappings in uniformly convex normed spaces, and then reused several years later in the generalised setting of uniformly convex hyperbolic spaces [30]. Similarly, quantitative variants of a useful lemma due to Xu [59], the first in [40], have found numerous applications in proof mining (typically for Halpern-type algorithms) which have been surveyed in [45]. A further family of related quantitative lemmas (see [43, Section 4.1]) have found applications in the very different contexts of contractive mappings [28, 47], set-valued operators [33] and jointly-nonexpansive families of mappings [52].
In the last few years, techniques from proof mining have been brought to bear on probability theory for the very first time [1, 42, 44], and this paper represents a new expansion into stochastic optimization. Given the similar significance of abstract lemmas for the convergence of stochastic algorithms, our main goal has been to provide a quantitative version of the most important of these – the Robbins-Siegmund theorem – and in this way open the gateway to providing new numerical results in the many concrete areas where that theorem is used.
We consider both the new connection between logic and stochastic optimization, together with the potential for future applications, to be the main achievements of the paper.
Methods from logic – notably proof interpretations based on Gödel’s Dialectica interpretation [15] – are important in both characterising the kind of quantitative information that might possible for a given theorem, extracting that information, and crucially in describing the underlying proof-theoretic phenomena which guarantee the extractability of such information. However, it is possible to present individual case studies without making explicit reference to those methods, and we adopt this approach in what follows, where details from logic and computability theory do not feature beyond this introduction.
Nevertheless, a high-level discussion of some relevant aspects of logic and computability is helpful in both outlining and motivating our main results, not least that convergence rates for the Robbins-Siegmund theorem – in a general and direct sense of the word – do not exist. We show that it is possible to provide alternative characterisations of convergence speed in terms of fluctuations or metastability (the latter in the sense of T. Tao [54]), and argue that for concrete applications, it is often extremely useful to have such information, again due to logical considerations, as we discuss below. It is precisely such an alternative characterisation in terms of metastability that we provide as our main result, and the remainder of this introduction is devoted to explaining this in more detail.
1.4. The impossibility of direct convergence rates
One of the simplest and yet most far-reaching facts from computable analysis – due to Specker [53] – is the existence of a monotone bounded sequence of rational numbers whose limit is not a computable real number. While the technical details of this construction involve concepts from the theory of computation (often formulated in terms of Turing machines), its consequences can be clearly stated without reference to these: Namely there is an algorithm for producing a monotone bounded sequence $x_{0},x_{1},x_{2},\ldots$ of rational numbers, whose speed of convergence to its limit $x^{\ast}$ is nevertheless impossible to characterise computationally, in the sense that there is no numerical method that on input $\varepsilon>0$ returns a number $N(\varepsilon)$ such that $|x_{n}-x^{\ast}|<\varepsilon$ for all $n\geq N(\varepsilon)$. In particular, this rules out any standard complexity bound (however slow) e.g. that $|x_{n}-x^{\ast}|\leq C/n$ or $C/\log n$ or $C/\log\log n$ and so on, as then $N(\varepsilon)$ could be readily computed (this general phenomenon that convergence can occur without computable rates is connected to the notion of arbitrary slow convergence that is more familiar in the optimization literature).
As a result, we see that there is no uniform way to manufacture rates of convergence for iterations satisfying Theorem 1.1 in terms of rates of convergence for $\sum_{i=0}^{\infty}\alpha_{i}<\infty$ and $\sum_{i=0}^{\infty}\gamma_{i}<\infty$, because in particular when $\alpha_{n}=\beta_{n}=\gamma_{n}=0$ we reobtain the monotone convergence theorem. We stress that this is not the case for all such recurrences: For example, if the nonnegative sequences $\{{x_{n}}\}$, $\{{\alpha_{n}}\}$ and $\{{\gamma_{n}}\}$ satisfy the closely related recurrence
$x_{n+1}\leq(1-\alpha_{n})x_{n}+\gamma_{n}$
for $\sum_{i=0}^{\infty}\alpha_{i}=\infty$ and $\sum_{i=0}^{\infty}\gamma_{i}<\infty$ (a situation that commonly arises when approximating fixpoints of nonexpansive operators [59]), then $x_{n}\to 0$ and one can indeed characterise the speed of convergence towards zero in terms of the divergence speed of $\sum_{i=0}^{\infty}\alpha_{i}=\infty$ and the convergence speed of $\sum_{i=0}^{\infty}\gamma_{i}<\infty$ (along with a bound on $x_{0}$), as made explicit in [40]. The key point here is that the condition $\sum_{i=0}^{\infty}\alpha_{i}=\infty$ imposes a strong enough descent on $\{{x_{n}}\}$ to guarantee fast convergence.
Noncomputability phenomena are directly inherited by both supermartingales and almost-supermartingales satisfying the premises of the Robbins-Siegmund theorem, for the simple reason that monotone decreasing sequences of reals are just special cases of real-valued supermartingales where $X_{n}(\omega)=x_{n}$ for all $\omega$. Thus, analogously to Theorem 1.1, there is no general way to construct direct, computable convergence rates for algorithms satisfying the premises of the Robbins-Siegmund theorem in terms of quantitative data on $\sum_{i=0}^{\infty}A_{i}<\infty$ and $\sum_{i=0}^{\infty}C_{i}<\infty$, either in the sense of rates of convergence for
$\lim_{N\to\infty}\mathbb{P}\left(\sup_{n\geq N}(|X_{n}-X^{\ast}|\geq\varepsilon)\right)=0$
or convergence rates for the individual trajectories $\{{X_{n}(\omega)}\}$. In line with this, most existing approaches to establishing convergence rates for algorithms reliant on the Robbins-Siegmund theorem (recent examples include [22, 41, 51]) focus on specific instances of the theorem where additional assumptions are imposed on the negative term $-B_{n}$ to ensure a strong descent and thus convergence rates. In this paper we take a very different approach, and instead adjust our notion of convergence speed, giving a quantitative version of the full Robbins-Siegmund theorem in terms of metastability. This is the first time such a quantitative result has been provided.
1.5. Our main result and its uses
The (Cauchy) convergence of a sequence of real numbers $\{{x_{n}}\}$ is equivalent to the following property, which (as set out in detail in Section 2.1 below) is derived from the usual formulation of convergence through simple logical reasoning and exchanging the order of the quantifiers:
For any $\varepsilon>0$ and $g:\mathbb{N}\to\mathbb{N}$ there exists some $n\in\mathbb{N}$ such that $|x_{i}-x_{j}|<\varepsilon$ for all $n\leq i,j\leq n+g(n)$.
In an analogous way, it can be shown that almost sure Cauchy convergence of some sequence of real-valued random variables $\{{X_{n}}\}$ can be reformulated as:
For any $\varepsilon,\lambda>0$ and $g:\mathbb{N}\to\mathbb{N}$ there exists some $n$ such that $|X_{i}-X_{j}|<\varepsilon$ for all $n\leq i,j\leq n+g(n)$ with probability $>1-\lambda$.
Both of the above properties can be viewed as “finitizations” of the original convergence property, in the sense that they replace the usual statement that the sequence eventually becomes permanently stable with metastability: stability on a finite range. The role that metastable convergence and other such finitizations play in different areas of mathematics has been discussed by Tao in the essay [54]. Here, finitary techniques along with finitized versions of theorems have proven particularly powerful in both ergodic theory and analytic number theory, notably the celebrated result of Green and Tao on arithmetic progressions of primes [16] (see also [55] for an example from ergodic theory where a convergence result is established by instead proving its metastable reformulation, and in which metastable variants of both ordinary and stochastic convergence are utilised).
Metastable convergence, along with the general process of finitizing mathematical statements, has a deep connection with logic, where the kind of “quantifier shuffling” involved in the finitization process is directly related to the proof interpretations mentioned in Section 1.3 that underlie the proof mining program. These techniques not only allow one to systematically derive the correct finitization of some specific theorem (see [14] for an interesting case study in this regard), but extract from the proof of that theorem explicit bounds.
The main result of this paper (Theorem 4.3) is the extraction of such bounds, specifically a bound on $n$ for metastable almost sure convergence above, from the proof of the Robbins-Siegmund theorem. We consider this to have several benefits.
Firstly, and in constract to existing work in this direction, we obtain a quantitative version of the full Robbins-Siegmund theorem without imposing any additional assumptions, which can moreover be viewed as a finitization of the theorem in the sense of Tao. This is of interest in its own right, particularly because the resulting metastable bounds are closely connected to bounds on fluctuations and upcrossings (a connection explored in more detail in [44]). These phenomena have long been studied in relation to martingales and ergodic theory (e.g. [5, 8, 17, 18, 19, 20, 21, 57]), and in this way our main result offers a novel extension of this line of research to encompass a very general class of almost-supermartingales.
More importantly, however, we envisage our main result being further developed and applied to obtain concrete numerical information from specific proofs that use the Robbins-Siegmund theorem. Indeed, an important benefit of the logical approach is that it works in a modular way, where informally speaking, as long as a theorem has been finitized in a suitably way, bounds for that finitary theorem can be “plugged in” directly to proofs that use it as a lemma, and in this way used as a component in a broader analysis.
Examples of this specific kind – where an abstract metastable convergence theorem has been used to obtain concrete numerical information in a specific setting – abound in the proof mining literature. For instance, a metastable variant of Theorem 1.1 is given in [29] and then applied to a Krasnoselskii-Mann algorithm to find a concrete bound on how many steps are needed until an approximate fixpoint of $T$ i.e. $\lVert{Tv_{n}-v_{n}}\rVert<\varepsilon$ is found. In Section 5 we provide a simple example exactly along these lines, in which our quantitative Robbins-Siegmund is used to bound the number of iterations of an abstract stochastic algorithm until an approximate solution is found. However, given the widespread use of the Robbins-Siegmund theorem across stochastic optimization, we anticipate that much deeper and far reaching applications are possible, and we outline several concrete ideas in this direction at the end of the paper.
We conclude this introduction by noting that, to date, proof mining program has had essentially no impact in probability or stochastic optimization. This is not due to any fundamental obstacle in the applicability of logical methods to stochastic algorithms, but a consequence of the fact that, until very recently, neither an appropriate quantitative understanding of martingales nor an abstract framework for explaining the relevant logical aspects of probability were available. These two steps have now been taken – by the authors and Pischke – in [44] and [42] respectively, thus opening the door to the application of logic in stochastic optimization. The present paper represents a first step in this direction.
1.6. Structure of the paper
In Section 2 we lay out the basic concepts and definitions that will be required, making precise several ideas that have only been informally discussed in this introduction. Then, as both a motivating example and an important lemma for our final section, we rederive the quantitative version of the nonstochastic Robbins-Siegmund theorem (Theorem 1.1), offering a different proof to [29] that more closely matches that of our main result. In Section 4 we state and prove our main result, before developing it further in Section 5, where we consider applications to abstract Robbins-Monro type algorithms. The proof of our main theorem relies heavily on previous work of the authors on martingales [44], though the results of that paper are much more general than those needed here. As such, we provide a streamlined proof of the specific convergence result we require in Appendix A, in an effort to make this paper completely self-contained. We reiterate that no knowledge of logic is required in any part of what follows, which relies on nothing more than some elementary analysis and the basics of martingale theory.
2. Quantitative convergence
This paper concerns quantitative convergence, so we begin by collecting together a number of preliminary definitions and lemmas on this theme. We first consider the simplified setting of nonstochastic convergence, as this helps motivate the main stochastic notions that follow. We use the notation
$[n;m]:=\{n,n+1,\ldots,m-1,m\}$
with the convention that $[n;m]=\emptyset$ if $m<n$. We freely use the logical quantifier symbols $\exists$,$\forall$ , ∀ throughout as they allow us to express several of our main definitions in a compact manner, and whenever these occur inside a probability measure they implicitly refer to an appropriate equivalent formulation using infinite unions or intersections: For example,
$\mathbb{P}\left(\forall\varepsilon>0\,\exists n\,(X_{n}<\varepsilon)\right)$
can be interpreted in a standard way as
$\mathbb{P}\left(\bigcap_{k\in\mathbb{N}}\bigcup_{n\in\mathbb{N}}(X_{n}<2^{-k})\right)$
In all such cases, the interpretation we choose will be clear.
2.1. Sequences of real numbers
Let $\{{x_{n}}\}$ be a sequence of reals, which is convergent iff it satisfies the Cauchy property:
(1) $\forall\varepsilon>0\,\exists n\,\forall i,j\geq n\left(|x_{i}-x_{j}|<\varepsilon\right)$
The most direct piece of quantitative information we could hope for is a function $f:(0,1)\to\mathbb{N}$ with the property that for each $\varepsilon>0$ we have $|x_{i}-x_{j}|<\varepsilon$ for all $i,j\geq f(\varepsilon)$. As outlined in Section 1.4, it is a basic fact of computable analysis that in general we cannot hope for $f$ to be computable, so with this in mind, we instead look consider alternative characterisations of convergence that in general does admit a computational interpretation. This motivates the following metastable reformulation of (1), which is nothing other that the property stated at the beginning of Section 1.5:
(2) $\forall\varepsilon>0\,\forall g:\mathbb{N}\to\mathbb{N}\,\exists n\,\forall i,j\in[n;n+g(n)]\left(|x_{i}-x_{j}|<\varepsilon\right)$
To see that (1) and (2) are equivalent, note that (1) clearly implies (2), and conversely, if (1) doesn’t hold then for some $\varepsilon>0$ there exists some $g$ such that
$\forall n\,\exists i,j\in[n;n+g(n)](|x_{i}-x_{j}|\geq\varepsilon)$
and taking the negation of this we obtain (2). A functional $\Phi:(0,1)\times\mathbb{N}^{\mathbb{N}}\to\mathbb{N}$ satisfying
$\forall\varepsilon>0\,\forall g:\mathbb{N}\to\mathbb{N}\,\exists n\leq\Phi(\varepsilon,g)\,\forall i,j\in[n;n+g(n)]\left(|x_{i}-x_{j}|<\varepsilon\right)$
is called a metastable rate of convergence (or just rate of metastability) for $\{{x_{n}}\}$. A detailed account of metastabile convergence and its connection to formal logical transformations can be found in [25] (see Chapter 2 in particular), while its significance in mainstream mathematics is discussed in [54]. In contrast to direct rates, metastable rates of convergence that are both computable and highly-uniform can often be obtained through analysing the respective convergence proofs, and indeed, the extraction of rates of metastability is a standard result in proof mining with numerous examples in recent years.
A very natural situation in which convergent sequences posses particularly simple metastable rates are those which for which an explicit bound on their $\varepsilon$-fluctuations can be given. We say that a sequence $\{{x_{n}}\}$ experiences $k$ $\varepsilon$-fluctuations if there exists
$i_{1}<j_{1}\leq i_{2}<j_{2}\leq\ldots\leq i_{k}<j_{k}\mbox{ with }|x_{i_{l}}-x_{j_{l}}|\geq\varepsilon$
for all $l=1,\ldots,k$. If $\phi:(0,1)\to\mathbb{R}$ is a function such that for any $\varepsilon>0$, $\{{x_{n}}\}$ has no more than $\phi(\varepsilon)$ $\varepsilon$-fluctuations, then the following is a metastable rate of convergence for $\{{x_{n}}\}$:
(3) $\Phi(\varepsilon,g):=\tilde{g}^{(\lceil\phi(\varepsilon)\rceil)}(0)$
where $\tilde{g}(n):=n+g(n)$ and $\tilde{g}^{(i)}$ denotes the $i$th iteration of $\tilde{g}$. This can be seen by observing that at least one of the intervals $[\tilde{g}^{(n)}(0);\tilde{g}^{(n+1)}(0)]$ for $n\leq\lceil\phi(\varepsilon)\rceil$ cannot contain any fluctuation. Indeed, as highlighted in [44, Section 2.3], a sequence enjoying a metastable rate of convergence of the simple form (3) is equivalent $\phi$ having the following property:
Definition 2.1.
Let $\{{x_{n}}\}$ be a sequence of real numbers. Any function $\phi:(0,1)\to\mathbb{R}$ such that for all $\varepsilon\in(0,1)$, for any increasing sequences
$a_{0}<b_{0}\leq a_{1}<b_{1}\leq\ldots$
there exists $n\leq\phi(\varepsilon)$ such that
$\forall i,j\in[a_{n};b_{n}]\,(|x_{i}-x_{j}|<\varepsilon)$
is called a learnable rate of convergence for $\{{x_{n}}\}$.
There are many situations in which convergent sequences possess learnable rates of convergence, and hence simple metastable rates of the form (3). The following result standard in proof mining (see in particular [25, Chapter 2]), and is also discussed by Tao in [54] where it is labelled the ‘finite convergence principle’:
Lemma 2.2.
Let $\{{x_{n}}\}$ be a nondecreasing sequence nonnegative reals such that $x_{n}<K$ for all $n\in\mathbb{N}$. Then a learnable rate of convergence for $\{{x_{n}}\}$ is given by $\phi(\varepsilon):=K/\varepsilon$, which in turn implies that a metastable rate of convergence is given by $\Phi(\varepsilon,g):=\tilde{g}^{\lceil K/\varepsilon\rceil}(0)$.
Proof.
Suppose for contradiction that for all $n\geq K/\varepsilon$ there exists some $i,j\in[a_{n};b_{n}]$ with $|x_{i}-x_{j}|=x_{j}-x_{i}\geq\varepsilon$. Then we have $K>x_{b_{n}}\geq(n+1)\varepsilon$ for all such $n$, but this is a contradiction for $n:=\lfloor K/\varepsilon\rfloor$. ∎
For example, in the special case of a convergent series $\sum_{i=0}^{\infty}\alpha_{i}<\infty$ where $\{{\alpha_{n}}\}$ is some sequence of nonnegative reals, if $K>0$ is any number satisfying $\sum_{i=0}^{\infty}\alpha_{i}<K$ then for any $a_{0}<b_{0}\leq a_{1}<b_{1}\leq\ldots$ then $\phi(\varepsilon):=K/\varepsilon$ is a learnable rate of convergence for $\{{\sum_{i=n}^{\infty}\alpha_{i}}\}$.
Metastable rates of the form (3) represent a simple instance of a more general kind of iterative metastable rate that arises from formulas that are effective learnability, a concept introduced and explored in [34], though for the purposes of this paper it is sufficient to restrict our attention to learnable rates of convergence in the simple sense of Definition 2.1 above. A notable advantage of having learnables rates (in our sense) is that they compose very easily under addition and multiplication of convergence sequences:
Lemma 2.3.
Suppose that $\{{x_{n}}\}$ and $\{{y_{n}}\}$ converge with learnable rates of convergence $\phi$ and $\psi$ respectively. Then
  1. (a)
    a learnable rate of convergence for $\{{x_{n}+y_{n}}\}$ is given by
    $\chi(\varepsilon):=\phi(\varepsilon/2)+\psi(\varepsilon/2)$
  2. (b)
    a learnable rate of convergence for $\{{x_{n}y_{n}}\}$ is given by
    $\chi(\varepsilon):=\phi(\varepsilon/2L)+\psi(\varepsilon/2K)$
    where $\sup_{n\in\mathbb{N}}|x_{n}|<K$ and $\sup_{n\in\mathbb{N}}|y_{n}|<L$.
We omit the proof, as it is just a simplified version of the proof of Lemma 2.10 below.
2.2. Sequences of real-valued random variables
We now introduce stochastic analogues of each of the concepts discussed above, where sequences of real numbers are replaced by sequences of random variables $\{{X_{n}}\}$ over some probability space $(\Omega,\mathcal{F},\mathbb{P})$, and we now talk about properties being true almost-surely. Informally speaking, we arrive at suitable quantitative notions of these properties by first using continuity properties of the probability function $\mathbb{P}$ to bring quantifiers from inside the measure (in the form of infinite unions and intersections) to the outside: See in particular [44, Section 3] for a detailed discussion of this process of ‘finitization’.
We first note that almost-sure Cauchy convergence of $\{{X_{n}}\}$ i.e.
$\mathbb{P}\left(\forall\varepsilon>0\,\exists n\,\forall i,j\geq n\left(|X_{i}-X_{j}|<\varepsilon\right)\right)=1$
is equivalent to the following metastable variant (see [2] where it is introduced and [44, Theorem 3.2] for the equivalence proof):
$\forall\lambda,\varepsilon>0\,\forall g:\mathbb{N}\to\mathbb{N}\,\exists n\,\left[\mathbb{P}\left(\exists i,j\in[n;n+g(n)]\left(|X_{i}-X_{j}|\geq\varepsilon\right)\right)<\lambda\right]$
This then leads to the following definition:
Definition 2.4 ([2, 3, 44]).
Let $\{{X_{n}}\}$ be a stochastic process. Any functional $\Phi:(0,1)\times(0,1)\times\mathbb{N}^{\mathbb{N}}\to\mathbb{N}$ such that for all $\lambda,\varepsilon\in(0,1)$ and $g:\mathbb{N}\to\mathbb{N}$ there exists $n\leq\Phi(\lambda,\varepsilon,g)$ satisfying
$\mathbb{P}(\exists i,j\in[n;n+g(n)](|X_{i}-X_{j}|\geq\varepsilon))<\lambda$
is called a metastable rate of uniform convergence for $\{{X_{n}}\}$.
Alternative metastable notions of almost sure convergence for stochastic processes have been proposed, including so-called metastable rates of pointwise convergence (cf. [2, 44]). However, pointwise rates do not directly feature in this paper: Metastable uniform convergence is stronger than pointwise convergence, and so the latter are only interesting in situations where rates for uniform convergence are difficult to achieve.
On the other hand, fluctuations for stochastic processes have been widely studied (for martingales, [20] is particularly relevant), and just as in the deterministic case, bounds on fluctuations give rise to simple iterative rates of metastability. Here the relevant stochastic analogue of learnable convergence is the following:
Definition 2.5 ([44]).
Let $\{{X_{n}}\}$ be a stochastic process. Any function $\phi:(0,1)\times(0,1)\to\mathbb{R}$ such that for all $\lambda,\varepsilon\in(0,1)$, for any increasing sequences
$a_{0}<b_{0}\leq a_{1}<b_{1}\leq\ldots$
there exists $n\leq\phi(\lambda,\varepsilon)$ such that
$\mathbb{P}\left(\exists i,j\in[a_{n};b_{n}]\,(|X_{i}-X_{j}|\geq\varepsilon)\right)<\lambda$
is called a learnable rate of uniform convergence for $\{{X_{n}}\}$.
The connection between learnable rates in the above sense and fluctuations is as follows:
Theorem 2.6 ([44]).
Let $\{{X_{n}}\}$ be a stochastic process and let $J_{\varepsilon}\{{X_{n}}\}$ denote the maximum number of $\varepsilon$-fluctuations experienced by $\{{X_{n}}\}$. Then
$\phi(\lambda,\varepsilon):=\frac{\mathbb{E}\left[J_{\varepsilon}\{{X_{n}}\}\right]}{\lambda}$
is a learnable rate of uniform convergence for $\{{X_{n}}\}$.
As such, learnable rates of uniform convergence are closely related to and generalise the property of a stochastic process having bounded fluctuations in mean. Entirely analogously to the deterministic case, whenever $\phi$ is a learnable rate of uniform convergence for $\{{X_{n}}\}$, then
$\Phi(\varepsilon,\lambda,g):=\tilde{g}^{(\lceil\phi(\lambda,\varepsilon)\rceil)}(0)$
is a metastable rate of uniform convergence for $\{{X_{n}}\}$.
In order to deal with infinite series along with sums and products of stochastic processes in a quantitative way, we must first give a quantitative interpretation to the statement that $\sup_{n\in\mathbb{N}}|X_{n}|<\infty$ almost surely. Again, bringing out the relevant quantifiers from within the probability measure function gives us the following notion (also discussed in [44, Section 4.1]):
Definition 2.7.
Let $\{{X_{n}}\}$ be a stochastic process. Any function $\rho:(0,1)\to\mathbb{R}$ such that
$\mathbb{P}\left(\sup_{n\in\mathbb{N}}|X_{n}|\geq\rho(\lambda)\right)<\lambda$
for all $\lambda\in(0,1)$ is called a modulus of uniform boundedness for $\{{X_{n}}\}$.
The following result on convergence for almost sure monotone sequences is important, and forms a stochastic analogue of Lemma 2.2 above.
Lemma 2.8.
Let $\{{X_{n}}\}$ be a nonnegative stochastic process such that $X_{n}\leq X_{n+1}$ almost surely for all $n\in\mathbb{N}$, and suppose that $\rho:(0,1)\to\mathbb{R}$ is a modulus of uniform boundedness for $\{{X_{n}}\}$. Then
$\phi(\lambda,\varepsilon):=\frac{2\cdot\rho(\lambda/2)}{\lambda\varepsilon}$
is a learnable rate of uniform convergence for $\{{X_{n}}\}$.
Proof.
Fix $\lambda,\varepsilon\in(0,1)$ and $a_{0}<b_{0}\leq a_{1}<b_{1}\leq\ldots$. Then for any $n\in\mathbb{N}$ we have
$\mathbb{P}\left(\exists i,j\in[a_{n};b_{n}]\,|X_{i}-X_{j}|\geq\varepsilon\right)$ $=\mathbb{P}\left(X_{b_{n}}-X_{a_{n}}\geq\varepsilon\right)$
$\leq\mathbb{P}\left(X_{b_{n}}-X_{a_{n}}\geq\varepsilon\cap\sup_{n\in\mathbb{N}}X_{n}<\rho(\lambda/2)\right)+\frac{\lambda}{2}$
Setting $Q_{n}:=X_{b_{n}}-X_{a_{n}}\geq\varepsilon$ and $R:=\sup_{n\in\mathbb{N}}X_{n}<\rho(\lambda/2)$ it suffices to show that there exists some $n\leq\phi(\lambda,\varepsilon)$ such that $\mathbb{P}(Q_{n}\cap R)<\lambda/2$. If this were not the case we would have
$\frac{\rho(\lambda/2)}{\varepsilon}<\frac{\lambda}{2}\left(\phi(\lambda,\varepsilon)+1\right)\leq\sum_{n=0}^{\phi(\lambda,\varepsilon)}\mathbb{P}(Q_{n}\cap R)=\mathbb{E}\left[I_{R}\sum_{n=0}^{\phi(\lambda,\varepsilon)}I_{Q_{n}}\right]$
But for $\omega\in R$ we can only have $\omega\in Q_{n}$ for strictly fewer than $\rho(\lambda/2)/\varepsilon$ distinct values of $n\in\mathbb{N}$, since whenever $\omega\in Q_{n_{1}}\cap\ldots\cap Q_{n_{k}}\cap R$ for $n_{1}<\ldots<n_{k}$ then
$\rho(\lambda/2)>X_{b_{n_{k}}}(\omega)\geq\sum_{j=1}^{k}\left(X_{b_{n_{j}}}(\omega)-X_{a_{n_{j}}}(\omega)\right)\geq k\varepsilon$
Therefore we obtain our contradiction by
$\mathbb{E}\left[I_{R}\sum_{n=0}^{\phi(\lambda,\varepsilon)}I_{Q_{n}}\right]\leq\mathbb{E}\left[I_{R}\cdot\frac{\rho(\lambda/2)}{\varepsilon}\right]\leq\frac{\rho(\lambda/2)}{\varepsilon}$
In particular, almost sure convergence of a series $\sum_{i=0}^{\infty}A_{i}$ where now the $A_{i}$ are nonnegative random variables can be represented quantitatively through a modulus of uniform boundedness $\rho$ i.e. a function satisfying
$\mathbb{P}\left(\sum_{i=0}^{\infty}A_{i}\geq\rho(\lambda)\right)<\lambda$
for all $\lambda\in(0,1)$, and this will be our preferred definition of quantitative convergence for infinite series. It is related to various other quantitative notions of convergence as follows:
Proposition 2.9.
Let $\{{A_{n}}\}$ be a sequence of nonnegative random variables.
  1. (a)
    If $\phi:(0,1)^{2}\to\mathbb{N}$ is a direct rate of convergence for $\sum_{i=0}^{\infty}A_{i}$ in the sense that
    $\mathbb{P}\left(\sum_{i=\phi(\lambda,\varepsilon)}^{\infty}A_{i}\geq\varepsilon\right)<\lambda$
    for all $\lambda,\varepsilon\in(0,1)$, and $a>0$ is such that $A_{n}\subseteq[0,a]$ for all $n\in\mathbb{N}$, then for any $\varepsilon\in(0,1)$
    $\rho_{\varepsilon}(\lambda):=a\cdot\phi(\lambda,\varepsilon)+\varepsilon$
    is a modulus of uniform boundedness for $\sum_{i=0}^{\infty}A_{i}$.
  2. (b)
    If $\rho:(0,1)\to\mathbb{N}$ is a modulus of uniform boundedness for $\sum_{i=0}^{\infty}A_{i}$ then
    $\psi(\lambda,\varepsilon):=\frac{2\cdot\rho(\lambda/2)}{\lambda\varepsilon}$
    is a learnable rate of uniform convergence for $\{{\sum_{i=0}^{n}A_{i}}\}$.
Proof.
For part (a) we observe that
$\mathbb{P}\left(\sum_{i=0}^{\infty}A_{i}\geq a\cdot\phi(\lambda,\varepsilon)+\varepsilon\right)$ $\leq\mathbb{P}\left(\sum_{i=0}^{\infty}A_{i}\geq\sum_{i=0}^{\phi(\lambda,\varepsilon)-1}A_{i}+\varepsilon\right)$
$=\mathbb{P}\left(\sum_{i=\phi(\lambda,\varepsilon)}^{\infty}A_{i}\geq\varepsilon\right)<\lambda$
Part (b) follows from Lemma 2.8 since $\{{\sum_{i=0}^{n}A_{i}}\}$ is monotone increasing. ∎
We now provide a stochastic version of Lemma 2.3 on quantitative convergence for sums and products, which will be exploited throughout the paper:
Lemma 2.10.
Suppose that $\{{X_{n}}\}$ and $\{{Y_{n}}\}$ converge with learnable rates of uniform convergence $\phi$ and $\psi$, respectively. Then
  1. (a)
    a learnable rate of uniform convergence for $\{{X_{n}+Y_{n}}\}$ is given by
    $\chi(\lambda,\varepsilon):=\phi(\lambda/2,\varepsilon/2)+\psi(\lambda/2,\varepsilon/2)$
  2. (b)
    a learnable rate of uniform convergence for $\{{X_{n}Y_{n}}\}$ is given by
    $\chi(\lambda,\varepsilon):=\phi\left(\frac{\lambda}{4},\frac{\varepsilon}{2\sigma(\lambda/4)}\right)+\psi\left(\frac{\lambda}{4},\frac{\varepsilon}{2\rho(\lambda/4)}\right)$
    where $\rho$ and $\sigma$ are moduli of uniform boundedness for $\{{X_{n}}\}$ and $\{{Y_{n}}\}$, respectively.
Proof.
We only prove (b), as part (a) is analogous but simpler (and in any case proven as Lemma 7.3 of [44]). Fix $\lambda,\varepsilon\in(0,1)$and define the events
$Q:=\sup_{n\in\mathbb{N}}|X_{n}|<\rho\left(\frac{\lambda}{4}\right)\ \ \ \mbox{and}\ \ \ R:=\sup_{n\in\mathbb{N}}|Y_{n}|<\sigma\left(\frac{\lambda}{4}\right)$
Take any $a_{0}<b_{0}\leq a_{1}<b_{1}\leq\ldots$. Then for any $n\in\mathbb{N}$ we have
$\mathbb{P}\left(\exists i,j\in[a_{n};b_{n}]\left(|X_{i}Y_{i}-X_{j}Y_{j}|\geq\varepsilon\right)\right)$
$\leq\mathbb{P}\left(\exists i,j\in[a_{n};b_{n}]\left(|X_{i}Y_{i}-X_{j}Y_{j}|\geq\varepsilon\right)\cap R\cap Q\right)+\frac{\lambda}{2}$
so it suffices to show that there exists some $n\leq\chi(\lambda,\varepsilon)$ such that
(4) $\mathbb{P}\left(\exists i,j\in[a_{n};b_{n}]\left(|X_{i}Y_{i}-X_{j}Y_{j}|\geq\varepsilon\right)\cap R\cap Q\right)<\frac{\lambda}{2}$
Take some $\omega$ in the set within the probability measure in (4) i.e. such that there exists $i:=i(\omega)$ and $j:=j(\omega)$ such that $|X_{i}(\omega)Y_{i}(\omega)-X_{j}(\omega)Y_{j}(\omega)|\geq\varepsilon$ and thus
$\varepsilon\leq\sigma(\lambda/4)|X_{i}(\omega)-X_{j}(\omega)|+\rho(\lambda/4)|Y_{i}(\omega)-Y_{j}(\omega)|$
which implies that $\sigma(\lambda/4)|X_{i}(\omega)-X_{j}(\omega)|\geq\varepsilon/2$ or $\rho(\lambda/4)|Y_{i}(\omega)-Y_{j}(\omega)|\geq\varepsilon/2$. Therefore if (4) fails for all $n\leq\chi(\lambda,\varepsilon)$, then we have either
$\mathbb{P}\left(\exists i,j\in[a_{n};b_{n}]\left(|X_{i}-X_{j}|\geq\frac{\varepsilon}{2\sigma(\lambda/4)}\right)\right)\geq\frac{\lambda}{4}$
or
$\mathbb{P}\left(\exists i,j\in[a_{n};b_{n}]\left(|Y_{i}-Y_{j}|\geq\frac{\varepsilon}{2\rho(\lambda/4)}\right)\right)\geq\frac{\lambda}{4}$
for each $n\leq\chi(\lambda,\varepsilon)$. As a consequence, there exists either a subsequence $a_{n_{0}}<b_{n_{0}}\leq a_{n_{1}}<b_{n_{1}}\leq\ldots$ such that
$\forall k\leq\phi\left(\frac{\lambda}{4},\frac{\varepsilon}{2\sigma(\lambda/4)}\right)\left[\mathbb{P}\left(\exists i,j\in[a_{n_{k}};b_{n_{k}}]\left(|X_{i}-X_{j}|\geq\frac{\varepsilon}{2\sigma(\lambda/4)}\right)\right)\geq\frac{\lambda}{4}\right]$
or a subsequence $a_{m_{0}}<b_{m_{0}}\leq a_{m_{1}}<b_{m_{1}}\leq\ldots$ such that
$\forall k\leq\psi\left(\frac{\lambda}{4},\frac{\varepsilon}{2\rho(\lambda/4)}\right)\left[\mathbb{P}\left(\exists i,j\in[a_{m_{k}};b_{m_{k}}]\left(|Y_{i}-Y_{j}|\geq\frac{\varepsilon}{2\rho(\lambda/4)}\right)\right)\geq\frac{\lambda}{4}\right]$
contradicting the defining properties of $\phi$ or $\psi$. ∎
We briefly conclude the section by making explicit how moduli of uniform boundedness combine under arithmetic operations, which is simple but essential for our main result:
Lemma 2.11.
Suppose that $\{{X_{n}}\}$ and $\{{Y_{n}}\}$ have moduli of uniform boundedness $\rho$ and $\sigma$, respectively. Then
  1. (a)
    a modulus of uniform boundedness for $\{{X_{n}+Y_{n}}\}$ is given by
    $\tau(\lambda):=\rho\left(\lambda/2\right)+\sigma\left(\lambda/2\right)$
  2. (b)
    a modulus of uniform boundedness for $\{{X_{n}Y_{n}}\}$ is given by
    $\tau(\lambda):=\rho\left(\lambda/2\right)\cdot\sigma\left(\lambda/2\right)$
Proof.
We just prove (b), as (a) is similar. Observe that for any $n,m\in\mathbb{N}$ we have
$\mathbb{P}\left(\sup_{n\in\mathbb{N}}|X_{n}Y_{n}|\geq nm\right)$ $\leq\mathbb{P}\left(\sup_{n\in\mathbb{N}}|X_{n}|\sup_{n\in\mathbb{N}}|Y_{n}|\geq nm\right)$
$\leq\mathbb{P}\left(\sup_{n\in\mathbb{N}}|X_{n}|\geq n\cup\sup_{n\in\mathbb{N}}|Y_{n}|\geq m\right)$
$\leq\mathbb{P}\left(\sup_{n\in\mathbb{N}}|X_{n}|\geq n\right)+\left(\sup_{n\in\mathbb{N}}|Y_{n}|\geq m\right)$
and so the result follows immediately. ∎
3. The nonstochastic Robbins-Siegmund theorem: A lemma of Qihou
Before presenting our quantitative Robbins-Siegmund theorem, we look briefly at the nonstochastic case, presented and discussed as Theorem 1.1 above, which serves as a motivating example where the structure of our main result and its proof are mirrored in simpler form, and connected to existing work in proof mining. The following result is proven in exactly the same way as Lemma 2.2, but in the stochastic setting, it will be replaced by a quantitative version of the Doob $L_{1}$ convergence theorem for supermartingales (rather than a variant of Lemma 2.8).
Lemma 3.1.
Let $\{{u_{n}}\}$ be a nonincreasing sequence of positive reals with $u_{0}<K$. Then $\phi(\varepsilon):=K/\varepsilon$ is a learnable rate of convergence for $\{{u_{n}}\}$.
We now give a quantitative version of Theorem 1.1, noting that the special case $\beta_{n}=0$ corresponds to a well-known lemma of Qihou [48, Lemma 2], which has already been analysed by Kohlenbach and Lambov in [29, Lemma 16]. We deliberately analyse a slightly different proof of the result, which forms a direct counterpart to the standard proof of the stochastic Robbins-Siegmund theorem. In particular, the reduction to a monotone sequences of positive reals exactly mirrors the reduction to a nonnegative supermartingale.
Theorem 3.2.
Let $\{{x_{n}}\}$, $\{{\alpha_{n}}\}$, $\{{\beta_{n}}\}$ and $\{{\gamma_{n}}\}$ be sequences of nonnegative reals with
$x_{n+1}\leq(1+\alpha_{n})x_{n}-\beta_{n}+\gamma_{n}$
for all $n\in\mathbb{N}$. Suppose that $K,L,M>0$ satisfy $x_{0}<K$, $\prod_{i=0}^{\infty}(1+\alpha_{i})<L$ and $\sum_{i=0}^{\infty}\gamma_{i}<M$. Then $\{{x_{n}}\}$ converges with learnable rate
$\phi(\varepsilon):=\frac{8L(K+M)}{\varepsilon}$
and moreover
$\sum_{i=0}^{\infty}\beta_{i}<L(K+M)$
Proof.
Define
$p_{n}:=\prod_{i=0}^{n-1}(1+\alpha_{i}),\ \ \ \tilde{x}_{n}:=\frac{x_{n}}{p_{n}},\ \ \ \tilde{\beta}_{n}:=\frac{\beta_{n}}{p_{n+1}},\ \ \ \tilde{\gamma}_{n}:=\frac{\gamma_{n}}{p_{n+1}}$
with $p_{0}=1$. For now we ignore $\{{\beta_{n}}\}$ and observe that
$\tilde{x}_{n+1}\leq\tilde{x}_{n}-\tilde{\beta}_{n}+\tilde{\gamma}_{n}\leq\tilde{x}_{n}+\tilde{\gamma}_{n}$
It then follows that
$u_{n}:=\tilde{x}_{n}-\sum_{i=0}^{n-1}\tilde{\gamma}_{n}$
with $u_{0}:=\tilde{x}_{0}$ is a nonincreasing sequence, and therefore $\{{u_{n}+M}\}$ is a nondecreasing, nonnegative sequence, and since $u_{0}=\tilde{x}_{0}\leq x_{0}<K$ we have $u_{0}+M<K+M$, and so by an application of Lemma 3.1 it follows that $\{{u_{n}+M}\}$ and hence $\{{u_{n}}\}$ has rate of uniform convergence $\phi_{1}(\varepsilon):=(K+M)/\varepsilon$. Similarly, since $\sum_{i=0}^{n-1}\tilde{\gamma}_{i}\leq\sum_{i=0}^{\infty}\gamma_{i}<M$, $\{{\sum_{i=0}^{n-1}\tilde{\gamma}_{i}}\}$ has rate of uniform convergence $\phi_{2}(\varepsilon):=M/\varepsilon$ (by Lemma 2.2). We now apply Lemma 2.3 repeatedly: First, $\{{\tilde{x}_{n}}\}$ has rate of uniform convergence
$\phi_{3}(\varepsilon):=\phi_{1}(\varepsilon/2)+\phi_{2}(\varepsilon/2)=\frac{2(K+M)}{\varepsilon}+\frac{2M}{\varepsilon}=\frac{2(K+2M)}{\varepsilon}$
and similarly since $x_{n}=\tilde{x}_{n}p_{n}$ and $\{{p_{n}}\}$ converges with learnable rate $\phi_{4}(\varepsilon):=L/\varepsilon$ (again by Lemma 2.2), a learnable rate for $\{{x_{n}}\}$ is given by
$\phi_{3}\left(\frac{\varepsilon}{2L}\right)+\phi_{4}\left(\frac{\varepsilon}{2K}\right)=\frac{4L(K+2M)}{\varepsilon}+\frac{2KL}{\varepsilon}\leq\frac{8L(K+M)}{\varepsilon}$
and the result for $\{{x_{n}}\}$ follows. For the second part of the theorem, note that
$\sum_{i=0}^{n}\tilde{\beta}_{i}=\tilde{x}_{0}-\tilde{x}_{n+1}+\sum_{i=0}^{n}\tilde{\gamma}_{i}\leq x_{0}+\sum_{i=0}^{n}\gamma_{i}<K+M$
and therefore
$\sum_{i=0}^{n}\beta_{i}=\sum_{i=0}^{n}\tilde{\beta}_{i}p_{n+1}<L\sum_{i=0}^{n}\tilde{\beta}_{i}<L(K+M)$
and the theorem is proved. ∎
It is instructive to compare our rate for $\{{x_{n}}\}$ with that obtained from Qihou’s lemma in [29, Lemma 16]. Their rate of metastability, where $L>0$ instead satisfies $\sum_{i=0}^{\infty}\alpha_{i}\leq L$ corresponds to the learnable rate
$\psi(\varepsilon):=\frac{4L(K+M)e^{L}+4M+(K+M)e^{L}}{\varepsilon}$
whereas ours, noting that if $L$ bounds the sum then $\prod_{i=0}^{\infty}(1+\alpha_{i})<e^{L}$, is just
$\phi(\varepsilon):=\frac{8(K+M)e^{L}}{\varepsilon}$
This difference reflects the more intricate argument used in [29] in comparison to our somewhat cruder strategy of reducing to a single monotone sequence then applying the sum and product rules Lemma 2.3 systematically. However, this strategy seems easier to extend to the stochastic setting, and is a direct reflection of standard proofs of the Robbins-Siegmund theorem.
4. The quantitative Robbins-Siegmund theorem
We now state and prove our quantitative Robbins-Siegmund theorem. We make use of the notions of quantitative convergence presented in Section 2, and the overall structure of both our main theorem and its proof mirror that of Theorem 3.2. Letting $\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\mathcal{F}_{2}\subseteq\ldots\mathcal{F}$ be some filtration of our underlying probability space $(\Omega,\mathcal{F},\mathbb{P})$, we recall that a supermartingale $\{{U_{n}}\}$ is some stochastic processes adapted to $\{{\mathcal{F}_{n}}\}$ such that $\mathbb{E}[|U_{n}|]<\infty$ and
$\mathbb{E}[U_{n+1}\mid\mathcal{F}_{n}]\leq U_{n}$
almost surely, for all $n\in\mathbb{N}$. We first need some facts about supermartingales. The following is standard:
Lemma 4.1 (Ville’s inequality).
Let $\{{U_{n}}\}$ be a nonnegative supermartingale. Then for any $a>0$ we have
$\mathbb{P}\left(\sup_{n\in\mathbb{N}}U_{n}\geq a\right)\leq\frac{\mathbb{E}[U_{0}]}{a}$
Alternatively, using our terminology, whenever $\mathbb{E}[U_{0}]<K$ then $\rho(\lambda):=K/\lambda$ is a modulus of uniform boundedness for $\{{U_{n}}\}$.
The next result on metastable rates of convergence for supermartingales falls out of the general framework of [44], though for the sake of completeness, a simplified proof of this special case is included in Appendix A. In the context of the proof that follows, it should be viewed as a stochastic analogue of Lemma 3.1.
Theorem 4.2 ([44]).
Let $\{{U_{n}}\}$ be a nonnegative supermartingale and suppose that $K>1$ is such that $\mathbb{E}[U_{0}]<K$. Then $\{{U_{n}}\}$ has learnable rate of uniform convergence given by:
$\phi(\lambda,\varepsilon):=c\left(\frac{K}{\lambda\varepsilon}\right)^{2}$
for a universal constant $c\leq 200$.
4.1. Main result
We are now ready for our main result, which represents a quantitative version of the Robbins-Siegmund theorem as formulated in the original paper [50, Theorem 1] or the more recent survey [13, Lemma 4.1] or indeed any number of standard texts on stochastic optimization (note that for simplicity we assume that all sequences involved are integrable). We assume that the moduli of uniform boundedness we need as an assumption are nonincreasing, which is a natural property and helps simplify the resulting rates.
Theorem 4.3 (Quantitative Robbins-Siegmund theorem).
Let $\{{X_{n}}\}$, $\{{A_{n}}\}$, $\{{B_{n}}\}$ and $\{{C_{n}}\}$ be nonnegative integrable stochastic processes adapted to some filtration $\{{\mathcal{F}_{n}}\}$ such that
$\mathbb{E}[X_{n+1}\mid\mathcal{F}_{n}]\leq(1+A_{n})X_{n}-B_{n}+C_{n}$
almost surely for all $n\in\mathbb{N}$. Suppose that $K>\mathbb{E}[X_{0}]$ and that $\rho,\sigma:(0,1)\to[1,\infty)$ are nonincreasing and satisfy
$\mathbb{P}\left(\prod_{i=0}^{\infty}(1+A_{i})\geq\rho(\lambda)\right)<\lambda\ \ \ \mbox{and}\ \ \ \mathbb{P}\left(\sum_{i=0}^{\infty}C_{i}\geq\sigma(\lambda)\right)<\lambda$
for all $\lambda\in(0,1)$. Then $\{{X_{n}}\}$ converges almost surely, with learnable rate of uniform convergence
$\phi(\lambda,\varepsilon):=\bar{c}\cdot\left(\frac{\rho\left(\frac{\lambda}{8}\right)\cdot\left(K+\sigma\left(\frac{\lambda}{16}\right)\right)}{\lambda\varepsilon}\right)^{2}$
where $\bar{c}>0$ is a constant that can be computed from that in Theorem 4.2, and $\sum_{i=0}^{\infty}B_{i}$ is finite almost surely, with modulus of uniform boundedness
$\chi(\lambda):=\frac{10\cdot\rho\left(\frac{\lambda}{2}\right)\cdot\left(K+\sigma\left(\frac{\lambda}{8}\right)\right)}{\lambda}$
Proof.
Our strategy is to analyse the standard proof of the result (as in [50]) making use of the quantitative lemmas we have already established. More precisely, we carry out the following steps:
  1. (1)
    Reduce $\{{X_{n}}\}$ via basic arithmetic operations to a stopped, positive supermartingale $\{{U_{n\wedge T_{x}}+x}\}$, rates of convergence and boundedness for which follow from Theorems 4.1 and 4.2.
  2. (2)
    Obtain rates for the unstopped supermartingale $\{{U_{n}}\}$.
  3. (3)
    Invert the reduction process to reobtain $\{{X_{n}}\}$, building up rates by using Lemmas 2.10 and 2.11.
First of all, analogously to the nonstochastic case, we define
$P_{n}:=\prod_{i=0}^{n-1}(1+A_{i}),\ \ \ \tilde{X}_{n}:=\frac{X_{n}}{P_{n}},\ \ \ \tilde{B}_{n}:=\frac{B_{n}}{P_{n+1}},\ \ \ \tilde{C}_{n}:=\frac{C_{n}}{P_{n+1}}$
with $P_{0}=1$. Since $\{{B_{n}}\}$ is nonnegative we have
$\mathbb{E}[X_{n+1}\mid\mathcal{F}_{n}]\leq(1+A_{n})X_{n}+C_{n}$
almost surely, and therefore defining
$U_{n}:=\tilde{X}_{n}-\sum_{i=0}^{n-1}\tilde{C}_{i}$
with $U_{0}:=\tilde{X}_{0}$ it follows that, almost surely,
$\mathbb{E}[U_{n+1}\mid\mathcal{F}_{n}]$ $=\mathbb{E}\left[\tilde{X}_{n+1}-\sum_{i=0}^{n}\tilde{C}_{i}\mid\mathcal{F}_{n}\right]$
$=\frac{\mathbb{E}\left[X_{n+1}\mid\mathcal{F}_{n}\right]}{P_{n+1}}-\sum_{i=0}^{n}\tilde{C}_{i}$
$\leq\frac{(1+A_{n})X_{n}+C_{n}}{P_{n+1}}-\sum_{i=0}^{n}\tilde{C}_{i}$
$=\tilde{X}_{n}-\sum_{i=0}^{n-1}\tilde{C}_{i}=U_{n}$
and thus $\{{U_{n}}\}$ is a supermartingale. Now, for $x>0$ define the stopping time $T_{x}\in\mathbb{N}\cup\{\infty\}$ by
$T_{x}:=\inf\left\{n\,:\,\sum_{i=0}^{n}\tilde{C}_{i}>x\right\}$
It is a standard fact that the stopped process $\{{U_{n\wedge T_{x}}}\}$ (where $n\wedge m:=\min\{n,m\}$) is also a supermartingale, since
$\mathbb{E}\left[U_{(n+1)\wedge T_{x}}\mid\mathcal{F}_{n}\right]=U_{T_{x}}I_{\left\{T_{x}\leq n\right\}}+\mathbb{E}\left[U_{n+1}I_{\left\{T_{x}>n\right\}}\mid\mathcal{F}_{n}\right]\leq U_{n\wedge T_{x}}$
and therefore $\{{U_{n\wedge T_{x}}+x}\}$ is a nonnegative supermartingale, since on $\{T_{x}<\infty\}$ we have
$U_{n\wedge T_{x}}+x=\tilde{X}_{n\wedge T_{x}}-\sum_{i=0}^{n\wedge T_{x}-1}\tilde{C}_{i}+x\geq X_{n\wedge T_{x}}-\sum_{i=0}^{T_{x}-1}\tilde{C}_{i}+x\geq X_{n\wedge T_{x}}\geq 0$
and on $\{T_{x}=\infty\}$ we have $\sum_{i=0}^{n-1}\tilde{C}_{i}\leq x$ for all $n\in\mathbb{N}$, and so also $U_{n\wedge T_{x}}+x\geq 0$. Noting that $\mathbb{E}[U_{0\wedge T_{x}}+x]=\mathbb{E}[\tilde{X}_{0}]+x<K+x$, by Theorem 4.2 a learnable rate of uniform convergence for $\{{U_{n\wedge T_{x}}+x}\}$ is given by
$\phi_{1}^{x}(\lambda,\varepsilon):=c\left(\frac{K+x}{\lambda\varepsilon}\right)^{2}$
where $c>0$ is the constant from that theorem. Clearly, this must then also be a learnable rate of uniform convergence for $\{{U_{n\wedge T_{x}}}\}$. Our next step is to find a rate for the unstopped sequence $\{{U_{n}}\}$. To this end, for $\lambda\in(0,1)$ define the event
$Q_{\lambda}:=\sum_{i=0}^{\infty}\tilde{C}_{i}<\sigma\left(\frac{\lambda}{2}\right)$
noting that since $\tilde{C}_{i}\leq C_{i}$ we have $\mathbb{P}(Q_{\lambda}^{c})<\lambda/2$. We now define
$\phi_{1}(\lambda,\varepsilon):=\phi^{\sigma(\lambda/2)}_{1}(\lambda/2,\varepsilon)=4c\left(\frac{K+\sigma(\lambda/2)}{\lambda\varepsilon}\right)^{2}$
Then for any $\lambda,\varepsilon\in(0,1)$ and $a_{0}<b_{0}\leq a_{1}<b_{1}\leq\ldots$ there exists some $n\leq\phi_{1}(\lambda,\varepsilon)$ such that
$\mathbb{P}\left(\exists i,j\in[a_{n};b_{n}]\left(|U_{i\wedge T_{\sigma(\lambda/2)}}-U_{j\wedge T_{\sigma(\lambda/2)}}|\geq\varepsilon\right)\right)<\lambda/2$
Now, we have
$\mathbb{P}\left(\exists i,j\in[a_{n};b_{n}]\left(|U_{i}-U_{j}|\geq\varepsilon\right)\right)\leq\mathbb{P}\left(\exists i,j\in[a_{n};b_{n}]\left(|U_{i}-U_{j}|\geq\varepsilon\right)\cap Q_{\lambda}\right)+\lambda/2$
and if $\omega\in Q_{\lambda}$ then $T_{\sigma(\lambda/2)}(\omega)=\infty$ and thus $U_{n\wedge T_{\sigma(\lambda/2)}(\omega)}(\omega)=U_{n}(\omega)$ for all $n\in\mathbb{N}$, and hence
$\mathbb{P}\left(\exists i,j\in[a_{n};b_{n}]\left(|U_{i}-U_{j}|\geq\varepsilon\right)\cap Q_{\lambda}\right)$
$\leq\mathbb{P}\left(\exists i,j\in[a_{n};b_{n}]\left(|U_{i\wedge T_{\sigma(\lambda/2)}}-U_{j\wedge T_{\sigma(\lambda/2)}}|\geq\varepsilon\right)\right)<\lambda/2$
from which it follows that $\phi_{1}$ is a learnable rate of uniform convergence for $\{{U_{n}}\}$. A similar argument allows us to exhibit a rate of uniform boundedness for $\{{U_{n}}\}$: By Ville’s inequality (Lemma 4.1), a modulus of uniform boundedness for the positive supermartingale $\{{U_{n\wedge T_{x}}+x}\}$ is given by
$\chi_{1}^{x}(\lambda):=\frac{K+x}{\lambda}$
It follows that
$\mathbb{P}\left(\sup_{n\in\mathbb{N}}|U_{n\wedge T_{x}}|\geq\frac{2(K+x)}{\lambda}\right)\leq\mathbb{P}\left(\sup_{n\in\mathbb{N}}\left(U_{n\wedge T_{x}}+x\right)\geq\frac{2(K+x)}{\lambda}\right)<\frac{\lambda}{2}$
where for the first inequality we note that for any $y>x$, if $|U_{n\wedge T_{x}(\omega)}(\omega)|\geq y$ then $|U_{n\wedge T_{x}(\omega)}(\omega)|=U_{n\wedge T_{x}(\omega)}(\omega)$ since $-U_{n\wedge T_{x}(\omega)}(\omega)\leq x$, and so in particular $U_{n\wedge T_{x}(\omega)}(\omega)+x\geq U_{n\wedge T_{x}(\omega)}(\omega)\geq y$, and since $2(K+x)/\lambda>x$ the inequality on suprema holds. Finally, we have
$\mathbb{P}\left(\sup_{n\in\mathbb{N}}|U_{n}|\geq\frac{2(K+\sigma(\lambda/2))}{\lambda}\right)$ $\leq\mathbb{P}\left(\left[\sup_{n\in\mathbb{N}}|U_{n}|\geq\frac{2(K+\sigma(\lambda/2))}{\lambda}\right]\cap Q_{\lambda}\right)+\frac{\lambda}{2}$
$\leq\mathbb{P}\left(\sup_{n\in\mathbb{N}}|U_{n\wedge T_{\sigma(\lambda/2)}}|\geq\frac{2(K+\sigma(\lambda/2))}{\lambda}\right)+\frac{\lambda}{2}<\lambda$
and so
$\chi_{1}(\lambda):=\frac{2(K+\sigma(\lambda/2))}{\lambda}$
is a modulus of uniform boundedness for $\{{U_{n}}\}$. We now conclude the proof using repeated applications of Lemmas 2.10 and 2.11.
First, since $\tilde{X}_{n}=U_{n}+\sum_{i=0}^{n-1}\tilde{C}_{i}$, and by $\tilde{C}_{i}\leq C_{i}$, $\sigma$ is also a modulus of uniform boundedness for $\{{\sum_{i=0}^{n-1}\tilde{C}_{n}}\}$, by Lemma 2.11 (a) a modulus of uniform boundedness for $\{{\tilde{X}_{n}}\}$ is given by
$\chi_{2}(\lambda):=\frac{4(K+\sigma(\lambda/4))}{\lambda}+\sigma(\lambda/2)$
and using Lemma 2.9 (b) in addition to Lemma 2.10 (a) a learnable rate of uniform convergence for $\{{\tilde{X}_{n}}\}$ is given by
$\phi_{2}(\lambda,\varepsilon):=64c\left(\frac{K+\sigma(\lambda/4)}{\lambda\varepsilon}\right)^{2}+\frac{8\sigma(\lambda/4)}{\lambda\varepsilon}$
For the final step, since $X_{n}=\tilde{X}_{n}P_{n}$ and $\rho$ is by definition a modulus of uniform boundedness for $\{{P_{n}}\}$, by Lemma 2.8 and monotonicity of $\{{P_{n}}\}$ a learnable rate of uniform convergence for the sequence is given by $2\rho(\lambda/2)/\lambda\varepsilon$, by Lemma 2.10 (b) a learnable rate of uniform convergence for $\{{X_{n}}\}$ is given by any bound on
$\phi_{2}\left(\frac{\lambda}{4},\frac{\varepsilon}{2\rho(\lambda/4)}\right)+\frac{16\cdot\chi_{2}(\lambda/4)\cdot\rho(\lambda/8)}{\lambda\varepsilon}$
The simplified bound in the statement of the theorem then follows in a crude way by bringing together terms, and using the assumptions $\lambda,\varepsilon\in(0,1)$ and that $\rho,\sigma$ are nonincreasing with $\rho(\lambda),\sigma(\lambda)\geq 1$.
To obtain the modulus of uniform boundedness on $\sum_{i=0}^{\infty}B_{i}$, we define
$V_{n}:=\tilde{X}_{n}-\sum_{i=0}^{n-1}(\tilde{C}_{i}-\tilde{B}_{i})=U_{n}+\sum_{i=0}^{n-1}\tilde{B}_{i}$
with $V_{0}:=\tilde{X}_{0}$. By an essentially identical argument to that for $\{{U_{n}}\}$, we can show that $\{{V_{n}}\}$ is a supermartingale, and defining the stopping time $T_{x}$ just as before and observing that because $\{{B_{n}}\}$ is nonnegative is it also the case that
$V_{n\wedge T_{x}}+x=U_{n\wedge T_{x}}+\sum_{i=0}^{n\wedge T_{x}-1}\tilde{B}_{i}+x\geq U_{n\wedge T_{x}}+x\geq 0$
on $\Omega$, it follows that $\{{V_{n\wedge T_{x}}+x}\}$ is a nonnegative supermartingale with $\mathbb{E}[V_{0\wedge T_{x}}+x]<K+x$ with modulus of uniform boundedness $\chi_{1}^{x}=(K+x)/\lambda$ as defined above. Then just as before, for $\omega\in Q_{\lambda}$ we have $T_{x}(\omega)=\infty$ and thus $V_{n\wedge T_{x}(\omega)}(\omega)=V_{n}(\omega)$, and so by an identical argument, $\chi_{1}$ as defined above is a modulus of uniform boundedness for $\{{V_{n}}\}$. We now apply Lemma 2.11 several times: Since
$\sum_{i=0}^{n-1}\tilde{B}_{i}=V_{n}+(-U_{n})\leq V_{n}+\sum_{i=0}^{n-1}C_{i}$
then
$\chi_{3}(\lambda):=\frac{5(K+\sigma(\lambda/4))}{\lambda}\geq\chi_{1}(\lambda/2)+\sigma(\lambda/2)$
is a modulus of uniform boundedness for $\{{\sum_{i=0}^{n-1}\tilde{B}_{i}}\}$. Finally, since
$\sum_{i=0}^{n-1}B_{i}=\sum_{i=0}^{n-1}\tilde{B}_{i}P_{i+1}\leq P_{n}\sum_{i=0}^{n-1}\tilde{B}_{i}$
a modulus of uniform boundedness for $\{{\sum_{i=0}^{n-1}B_{i}}\}$ is given by Lemma 2.11 as
$\chi(\lambda):=\rho(\lambda/2)\cdot\chi_{3}(\lambda/2)$
Making this explicit yields the second part of the theorem. ∎
We conclude by noting that, as a useful byproduct of the proof above, we obtain a modulus of uniform boundedness for $\{{X_{n}}\}$.
Corollary 4.4.
Under the assumptions of Theorem 4.3, a modulus of uniform boundedness for $\{{X_{n}}\}$ is given by
$\tau(\lambda):=\frac{9\left(K+\sigma\left(\tfrac{\lambda}{8}\right)\right)\cdot\rho\left(\tfrac{\lambda}{2}\right)}{\lambda}$
Proof.
Letting $\chi_{2}$ and $\rho$ be the moduli of uniform boundedness for $\{{\tilde{X}_{n}}\}$ and $\{{P_{n}}\}$ as defined in the proof of Theorem 4.3, by Lemma 2.11, $\chi_{2}(\lambda/2)\cdot\rho(\lambda/2)$ is a modulus of uniform boundedness for $X_{n}=\tilde{X}_{n}P_{n}$, and the result follows by substituting in the definition of $\chi_{2}$ and observing that $\tau(\lambda)\geq\chi_{2}(\lambda/2)\cdot\rho(\lambda/2)$. ∎
4.2. Comparison with recent quantitative results on stochastic gradient descent
Our main theorem, with its focus on logical methods and metastable convergence, is completely different in approach to most existing methods aimed at producing quantitative results for stochastic algorithms based on the Robbins-Siegmund theorem, which, as mentioned in the introduction, tend to focus on special cases of the theorem where direct rates are possible. Indeed, our quantitative results have more in common with fluctuation bounds for martingales, whose relationship with metastable convergence is discussed in detail in [44].
Moreover, we claim that our approach complements and in some cases even enhances existing convergence rates. To illustrate this, we take one of the most recent results in this direction: Theorem 2 of [22] (a paper that in turn builds on [41, 51]). Informally speaking, this considers an instance of the Robbins-Siegmund theorem with strong descent assumptions (where convergence of $\{{X_{n}}\}$ can then in turn be strengthened to $X_{n}\to 0$), and states that for any $\mu\in(0,1]$, whenever the parameters of the main recurrence converge or diverge sufficiently fast (with respect to $\mu$) then $X_{n}$ is $o(n^{-\mu})$ almost surely. This is proven through another instance of Robbins-Siegmund, in which a substitution is made so that the theorem is now applied to $\tilde{X}_{n}:=X_{n}n^{\mu}$ and used to show that $\tilde{X}_{n}\to 0$ almost surely.
The work in our paper can be used to augment results of this kind with additional quantitative information. For example, given suitable moduli representing convergence/divergence of the parameters, from Corollary 4.4 we would obtain a modulus of uniform boundedness $\rho$ for $\tilde{X}_{n}=X_{n}n^{\mu}$ i.e.
$\mathbb{P}\left(\sup_{n\in\mathbb{N}}X_{n}n^{\mu}\geq\rho(\lambda)\right)<\lambda$
or alternatively,
$\mathbb{P}\left(\forall n\left(X_{n}<\rho(\lambda)n^{-\mu}\right)\right)>1-\lambda$
In other words, while from $X_{n}$ being $o(n^{-\mu})$ almost surely we obviously have that $X_{n}$ is $O(n^{-\mu})$ almost surely, our modulus of uniform boundedness allows us to compute, for any $\lambda>0$, an explicit constant for $X_{n}$ being $O(n^{-\mu})$, uniform over different trajectories, that is valid with probably $1-\lambda$. Similarly, Corollary 5.4 below gives us an explicit uniform bound for finding some point such that $X_{n}<\varepsilon n^{-\mu}$ with probably $1-\lambda$, and further results along these lines are possible, all of which augment the results of [22] with additional quantitative information.
As such, we propose that our approach through metastability represents a complementary technique in comparison to work that focuses on providing rates of convergence for instances of the Robbins-Siegmund theorem, which through the existence of Specker sequences (as discussed in Section 1.4) can in any case never apply to the theorem in its general form, as our results do. On the other hand, we now show that in special cases, our main theorem can be used to provide concrete quantitative information under very weak and abstract conditions.
5. Discussion: Convergence of abstract stochastic algorithms
The purpose of this final section is to illustrate how the different components of Theorem 4.3 can be combined with slightly stronger assumptions to produce new abstract convergence results that resemble specific strategies for using the Robbins-Siegmund theorem. Much is possible in this general direction, and we give several specific ideas for future research in Section 6. For the purpose of the present paper, our aim is simply to discuss one of the most basic and widely used strategies for applying the Robbins-Siegmund theorem, where both parts of the theorem are used together to prove the stronger statement $X_{n}\to 0$. Typically, $X_{n}$ is related to $\lVert{x_{n}-\theta}\rVert$ for some stochastic algorithm $\{{x_{n}}\}$ intended to approximate an object $\theta$, where from $X_{n}\to 0$ we are then able to infer that $x_{n}\to\theta$. Classic examples of algorithms whose convergence can be proven using this strategy include the Robbins-Monro [49] and Kiefer-Wolfowitz [23] schemes, though in this paper we work on a much higher level of abstraction.
To represent such an approach in the abstract, throughout this section we consider the special case of Theorem 4.3 where $B_{n}:=u_{n}V_{n}$ for $\{{V_{n}}\}$ some nonnegative integrable stochastic process and $\{{u_{n}}\}$ a sequence of nonnegative reals – in other words, we assume that
(5) $\mathbb{E}[X_{n+1}\mid\mathcal{F}_{n}]\leq(1+A_{n})X_{n}-u_{n}V_{n}+C_{n}$
almost surely for all $n\in\mathbb{N}$ and that the other conditions of Theorem 4.3 are satisfied with respect to some fixed $K>\mathbb{E}[X_{0}]$ and moduli of uniform boundedness $\rho,\sigma$ for $\prod(1+A_{i}),\sum C_{i}<\infty$ respectively. We denote the resulting learnable rate of convergence for $\{{X_{n}}\}$ by $\phi$ and modulus of uniform boundedness for $\sum u_{i}V_{i}$ by $\chi$, both as defined in Theorem 4.3, and the modulus of uniform boundedness for $\{{X_{n}}\}$ by $\tau$, as defined in Corollary 4.4.
Intuitively, the $\{{u_{n}}\}$ perform the role of step-sizes, and for that reason we also require that $\sum u_{i}=\infty$. Because we are concerned with quantitative results, we require in addition a so-called rate of divergence for this series, which here will be a function $r:\mathbb{N}\times(0,\infty)\to\mathbb{N}$ such that
$\forall n,x\left(\sum_{i=n}^{n+r(n,x)}u_{i}\geq x\right)$
An immediate corollary of the Robbins-Siegmund theorem in this case is that
$\liminf_{n\to\infty}V_{n}=0$
We make this property quantitative by introducing a corresponding modulus:
Definition 5.1.
A function $\Phi:(0,1)\times(0,1)\times\mathbb{N}\to\mathbb{N}$ is a (stochastic) liminf-modulus for a sequence of random variables $\{{V_{n}}\}$ if for all $\lambda,\varepsilon>0$ and $n\in\mathbb{N}$:
$\mathbb{P}\left(\forall k\in[n;n+\Phi(\lambda,\varepsilon,n)](V_{k}\geq\varepsilon)\right)<\lambda$
The following is now straightforward:
Corollary 5.2.
Under the above conditions, a liminf-modulus for $\{{V_{n}}\}$ is given by
$\Phi(\lambda,\varepsilon,n):=r\left(n,\frac{\chi(\lambda)}{\varepsilon}\right)$
Proof.
Fix $\lambda,\varepsilon>0$ and $n\in\mathbb{N}$, and take some $\omega\in\Omega$. Suppose that $V_{i}(\omega)\geq\varepsilon$ for all $k\in[n;n+\Phi(n,\varepsilon)]$. Then
$\chi(\lambda)\leq\varepsilon\sum_{i=n}^{n+\Phi(\lambda,\varepsilon,n)}u_{i}\leq\sum_{i=n}^{n+\Phi(\lambda,\varepsilon,n)}u_{i}V_{i}(\omega)\leq\sum_{i=0}^{\infty}u_{i}V_{i}(\omega)$
and therefore we have shown that
$\mathbb{P}\left(\forall k\in[n;n+\Phi(\lambda,\varepsilon,n)](V_{k}\geq\varepsilon)\right)\leq\mathbb{P}\left(\sum_{i=0}^{\infty}u_{i}V_{i}\geq\chi(\lambda)\right)<\lambda$
which completes the proof. ∎
In the particular situations which we seek to present in the abstract, there would be a concrete relationship between $\{{V_{n}}\}$ and $\{{X_{n}}\}$ which allows us to infer $\liminf_{n\to\infty}X_{n}$ from $\liminf_{n\to\infty}V_{n}$. For example, in the well application of the Robbins-Siegmund theorem to prove convergence of the Robbins-Monro algorithm, following the presentation of [39] (cf. page 393) the main recurrence has the form (5) for $X_{n}:=(x_{n+1}-\theta)^{2}$ and $V_{n}:=M(x_{n+1})(x_{n+1}-\theta)$, where $M$ is a measurable function satisfying
$\inf_{\varepsilon\leq|x-\theta|\leq\varepsilon^{-1}}\left\{M(x)(x-\theta)\right\}>0$
Similarly, the more recent quantitative analysis of stochastic gradient descent [22, Theorem 1] asserts the existence of a function $\eta:(0,\infty)\to(0,\infty)$ satisfying
$V_{n}(\omega)\geq\eta(X_{n}(\omega))$
where $\eta$ has the additional property that for any $0<\varepsilon\leq M$
$\inf_{\varepsilon\leq r\leq K}\eta(r)>0$
We can give such relationships a general quantitative form, which also facilitates a quantitative route between the corresponding liminf-moduli, in the following way:
Lemma 5.3.
Suppose that $\delta:(0,\infty)\times(0,\infty)\to(0,\infty)$ satisfies
(6) $\varepsilon\leq X_{n}(\omega)\leq K\implies V_{n}(\omega)\geq\delta(\varepsilon,K)$
for all $0<\varepsilon\leq K$ and $\omega\in\Omega$. Then a liminf-modulus for $\{{X_{n}}\}$ is given by
$\Psi(\lambda,\varepsilon,n):=\Phi\left(\lambda/2,\delta\left(\varepsilon,\tau(\lambda/2)\right),n\right)$
where $\Phi$ is any liminf-modulus for $\{{V_{n}}\}$, and so in particular that defined in Corollary 5.2.
Proof.
Fix $\lambda,\varepsilon>0$ and $n$ and define the events
$A:=\sup_{k\in\mathbb{N}}X_{k}<\tau(\lambda/2)\ \ \ \mbox{and}\ \ \ B:=\forall k\in[n;n+\Psi(\lambda,\varepsilon,n)](X_{k}\geq\varepsilon)$
where we recall that $\tau$ is a modulus of uniform boundedness for $\{{X_{n}}\}$. Take $\omega\in A\cap B$. Then for any $k\in[n;n+\Psi(\lambda,\varepsilon,n)]$ we have $\varepsilon\leq X_{k}(\omega)<\tau(\lambda/2)$ and thus
$V_{k}(\omega)\geq\delta\left(\varepsilon,\tau(\lambda/2)\right)$
We have therefore shown that
$\mathbb{P}(A\cap B)\leq\mathbb{P}(\forall k\in[n;n+\Psi(\lambda,\varepsilon,n)](V_{k}\geq\delta\left(\varepsilon,\tau(\lambda/2)\right))<\lambda/2$
and therefore
$\mathbb{P}(B)\leq\mathbb{P}(A\cap B)+\mathbb{P}(A^{c})<\lambda/2+\lambda/2=\lambda$
and we are done. ∎
To give a concrete example of a modulus $\delta$ as in Theorem 5.3, in the case of the Robbins-Monro algorithm discussed above, for any function $\mu:(0,\infty)\to(0,\infty)$ satisfying
$\inf_{\varepsilon\leq|x-\theta|\leq\varepsilon^{-1}}\left\{M(x)(x-\theta)\right\}\geq\mu(\varepsilon)$
the function
$\delta(\varepsilon,K):=\mu\left(\sqrt{\min\left\{\varepsilon,1/K\right\}}\right)$
would satisfy (6).
Putting everything together now gives us the following main convergence result, which can be seen as a special case of Theorem 4.3 for abstract Robbins-Monro-like algorithms.
Theorem 5.4.
Suppose that $\{{X_{n}}\},\{{V_{n}}\},\{{A_{n}}\},\{{C_{n}}\}$ and $\{{u_{n}}\}$ satisfy (5), that $K>\mathbb{E}[X_{0}]$ and that $\rho,\sigma$ are nonincreasing moduli of uniform boundedness for $\prod(1+A_{i}),\sum C_{i}<\infty$ respectively. Suppose that $\sum u_{i}=\infty$ with rate $r$, and that (6) is satisfied for some function $\delta$. Then $X_{n}\to 0$ almost surely, and moreover for any $\lambda,\varepsilon>0$ and $g:\mathbb{N}\to\mathbb{N}$ there exists some $n\leq\Gamma(\lambda,\varepsilon,g)$ such that
$\mathbb{P}\left(\exists k\in[n;n+g(n)](X_{k}\geq\varepsilon)\right)<\lambda$
where
$\Gamma(\lambda,\varepsilon,g):=\tilde{f}^{(\lceil\phi(\lambda/2,\varepsilon/2)\rceil)}(0)\ \ \ \mbox{for}\ \ \ f(j):=\max\{g(j),\Psi(\lambda/2,\varepsilon/2,j)\}$
for $\tilde{f}(j):=j+f(j)$, and where $\Psi$ is defined as in Lemma 5.3 for $\Phi$ as defined in Corollary 5.2 and $\phi,\chi$ and $\tau$ as defined in Theorem 4.3 and Corollary 4.4.
Proof.
Recall that $\phi$ is a learnable rate of uniform convergence for $\{{X_{n}}\}$, and so now fixing $\lambda,\varepsilon>0$, there exists some $n\leq\tilde{f}^{(\lceil\phi(\lambda/2,\varepsilon/2)\rceil)}(0)$ such that
$\mathbb{P}\left(\exists i,j\in[n;n+f(n)](|X_{i}-X_{j}|\geq\varepsilon/2)\right)<\lambda/2$
Let $D$ be the event inside the probability above. For this particular $n$, since $\Psi$ is a liminf modulus for $\{{X_{n}}\}$ we also have
$\mathbb{P}\left(\forall k\in[n;n+\Psi(\lambda/2,\varepsilon/2,n)](X_{k}\geq\varepsilon/2)\right)<\lambda/2$
Let $E$ be the event inside this probability. Fix $\omega\in\Omega$ and suppose that there exists some $k(\omega)\in[n;n+g(n)]\subseteq[n;n+f(n)]$ such that $X_{k(\omega)}(\omega)\geq\varepsilon$. Either $\omega\in E$, or there exists some $j(\omega)\in[n;n+\Psi(\lambda/2,\varepsilon/2,n)]\subseteq[n;n+f(n)]$ such that $X_{j(\omega)}(\omega)<\varepsilon/2$. But this then implies that $|X_{j(\omega)}(\omega)-X_{k(\omega)}(\omega)|\geq\varepsilon/2$, and so $\omega\in D$. Therefore
$\mathbb{P}\left(\exists k\in[n;n+g(n)](X_{k}\geq\varepsilon)\right)\leq\mathbb{P}(D\cup E)\leq\mathbb{P}(D)+\mathbb{P}(E)<\lambda$
and the theorem is proved. ∎
As expected, the conclusion of Theorem 5.4 does not provide a direct rate of convergence for $X_{n}\to 0$, only a metastable version. However, it should be noted that even the weaker liminf-modulus $\Psi$ for $\{{X_{n}}\}$ constructed in Lemma 5.3 provides relevant numerical information on the convergence speed of the underlying algorithm: In particular, defining $f(\lambda,\varepsilon):=\Psi(\lambda,\varepsilon,0)$ we have
$\mathbb{P}\left(\exists k\leq f(\lambda,\varepsilon)(X_{k}<\varepsilon)\right)>1-\lambda$
In a typical concrete scenario where we might have e.g. $X_{n}:=\lVert{x_{n}-\theta}\rVert$ for some stochastic algorithm $\{{x_{n}}\}$ converging to a solution $\theta$, we have therefore shown that up to any desired degree of accuracy $\lambda>0$ there exists a set $A_{\lambda}$ with $\mathbb{P}(A_{\lambda}^{c})<\lambda$ along with a function $\psi_{\lambda}:=f(\lambda,\cdot)$ such that for any $\omega\in A_{\lambda}$, the trajectory $\{{x_{n}(\omega)}\}$ produces an approximate $\varepsilon$-solution $x_{k}(\omega)$ – in the sense that $\lVert{x_{k}(\omega)-\theta}\rVert<\varepsilon$ – within the first $\psi_{\lambda}(\varepsilon)$ iterates. In general, quantitative information of this kind has a clear practical significance.
To give a concrete example of the general constructions above, if the step sizes $u_{i}:=u$ are constant, then we can define our rate of divergence $r(n,x):=\lceil x/u\rceil$, and so
$\Phi(\lambda,\varepsilon,n)=\left\lceil\frac{\chi(\lambda)}{u\cdot\varepsilon}\right\rceil=\left\lceil\frac{10\cdot\rho\left(\frac{\lambda}{2}\right)\cdot\left(K+\sigma\left(\frac{\lambda}{8}\right)\right)}{u\cdot\lambda\varepsilon}\right\rceil$
which in turn implies that
$f(\lambda,\varepsilon):=\Psi(\lambda,\varepsilon,0)=\left\lceil\frac{20\cdot\rho\left(\frac{\lambda}{4}\right)\cdot\left(K+\sigma\left(\frac{\lambda}{16}\right)\right)}{u\cdot\lambda\cdot\delta(\varepsilon,\tau(\lambda/2))}\right\rceil$
Further assumptions then simplify this function: If, for example, we have $\prod_{i=0}^{\infty}(1+A_{i})<L$ and $\sum_{i=0}^{\infty}C_{i}<M$ almost surely, then the corresponding moduli of uniform boundedness are given by constant functions $\rho(\lambda):=L$ and $\sigma(\lambda):=M$. If, furthermore, condition (6) is strengthened to
$X_{n}(\omega)\geq\varepsilon\implies V_{n}(\omega)\geq\delta(\varepsilon)$
it follows that
$f(\lambda,\varepsilon)\leq\frac{d}{\lambda\cdot\delta(\varepsilon)}$
for a suitable constant $d$ definable from $K,L,M$ and $u$, and so in particular a bound on the search for approximate solutions is given by
$\mathbb{P}\left(\exists k\leq\frac{d}{\lambda\cdot\delta(\varepsilon)}\,(X_{k}<\varepsilon)\right)>1-\lambda$
This demonstrates that the general quantitative results in the main part of this paper can, under additional assumptions, be used to obtain numerical information that is both simple and highly uniform, which would in turn be inherited by any stochastic algorithms whose convergence reduces to that particular instance of the Robbins-Siegmund theorem.
Indeed, while we consider the main result of this paper to be of interest in its own right as a quantitative, metastable version of a major theorem in stochastic approximation, our work is primarily intended as a basis for advancing logic-based methods – already highly successful in nonstochastic optimization – into the realm of stochastic algorithms. While we have hinted at this in the current section from an abstract perspective, we conclude by outlining several directions in which the abstract results presented here can be extended and applied to concrete algorithms.
6. Future work
6.1. Stochastic Fejér monotonicity
One of our main intended applications for Theorem 4.3 is providing quantitative information for stochastic quasi-Fejér monotone sequences – studied in a simplified form in the context of generalized gradient descent methods in [12] and reintroduced in a more general setting in [6]. Following the presentation of [6], we say that a stochastic process $\{{x_{n}}\}$ in some Hilbert space $X$ is quasi-Fejér monotone with respect to the set $Z$ if for all $z\in Z$ there exist random sequences $\{{A_{n}(z)}\}$, $\{{B_{n}(z)}\}$ and $\{{C_{n}(z)}\}$ adapted to $\{{\mathcal{F}_{n}}\}$ such that
$\mathbb{E}[\phi(\lVert{x_{n+1}-z}\rVert)\mid\mathcal{F}_{n}]\leq(1+A_{n}(z))\phi(\lVert{x_{n}-z}\rVert)-B_{n}(z)+C_{n}(z)$
for all $n\in\mathbb{N}$, where $\phi:[0,\infty)\to[0,\infty)$ is a strictly increasing function with $\lim_{t\to\infty}\phi(t)=\infty$ and $\prod(1+A_{n}(z)),\sum C_{n}(z)<\infty$ almost surely. In [6], the Robbins-Siegmund theorem plays a central role in establishing the convergence of $\{{x_{n}}\}$ under additional assumptions, which can then in turn be used to prove convergence of a number of concrete iterative algorithms in Hilbert spaces.
In upcoming work with Pischke, we analyse the main proof of [6] to obtain a general quantitative convergence result – valid in a much more broader setting – for stochastic quasi-Fejér monotone sequences, building on a precedent already well established within proof mining for nonstochastic Fejér monotonicity [31, 32, 46]. Our quantitative Robbins-Siegmund theorem is crucial for this result, where here we even require a strengthening of Theorem 4.3 in which $\{{X_{n}}\}$ is now a finite supermartingale, which in turn involves the development of a finitary theory of martingales. In this way, our finitisation of the Robbins-Siegmund theorem can be further expanded, while at the same time providing numerical results that are immediately applicable to concrete stochastic algorithms, including, for example, the block-coordinate fixed point algorithms of [6], for which no quantitative convergence information currently exists.
6.2. Convergence rates for classic algorithms
An obvious application of the quantitative results of this paper is to obtain numerical information for stochastic algorithms for which rates of convergence are not known. However, in Section 5 of this paper we briefly discussed ways in which our main theorems could be used to produce abstract versions of classic schemes, for which there exists an extensive literature on convergence rates (see e.g. [10, 38]). Therefore an interesting open question here is whether our logic-based approach could be used to find numerical information for generalised versions of those schemes that augments existing knowledge in a useful way. In particular, we claim that under further strengthenings of the conditions in Section 5, where instead of a uniform bound for $\sum_{i=0}^{\infty}C_{i}$ we have a rate of convergence for $\sum_{i=0}^{\infty}\mathbb{E}[C_{i}]$, we can use concentration inequalities such as Ville’s inequality to upgrade our rates for approximate solutions to full convergence rates. This would then in turn then allow us to provide general convergence rates for abstract Robbins-Monro type algorithms, which might be formulated in terms of abstract moduli such as (6) rather than e.g. explicit properties on some regression function whose root is to be approximated. This is also ongoing work with the authors and Pischke.
A particularly interesting case study is represented by Dvoretsky’s approximation theorem [11]. This is less general than the Robbins-Siegmund theorem but nevertheless encompasses the Robbins–Monro and Kiefer–Wolfowitz schemes as special cases. It is well known that Dvoretsky’s theorem is provable as a special case of Robbins-Siegmund theorem [50], and so identifying the quantitative version of this theorem that results from applying Theorem 4.3 would be illuminating. An important point to make here is that in the case of Dvoretsky’s theorem, more elementary proofs that do not make use of martingale theory are available, notably that of Derman and Sacks [7]. In particular, analysing the proof of Derman and Sacks would likely give us a different quantitative result, and it is an open question whether this would result in better rates than those obtained through the use of supermartingales [^1] We are grateful to Rob Arthan and Paulo Oliva, who are currently working on such a quantitative analysis of the Derman-Sacks proof, for raising this point., though our quantitative Robbins-Siegmund theorem is much more broadly applicable.
6.3. A toolkit for analysing general recursive inequalities
Finally, we mention that while the Robbins-Siegmund theorem may be the most well-known manifestations of almost-supermartingale convergence in stochastic approximation, there exist numerous variants, by which we mean abstract convergence results of the general form that from some supermartingale-like property
$\mathbb{E}[X_{n+1}\mid\mathcal{F}_{n}]\leq\phi_{n}(X_{n})$
for a particular sequence of functions $\phi_{n}$, then something can be said about the convergence of $\{{X_{n}}\}$. Several examples of this kind can be found in e.g. [10] or the recent survey paper [13] (see Section 4 in particular), and we hope that the many lemmas and auxiliary results of this paper, particularly those in Section 2 and Section 5, will form a useful toolkit for establishing quantitative versions of alternative convergence results in the stochastic setting.
Related to this point, the authors have previously discussed the merits of using abstract recursive inequalities as the a basic organisational structure of a library of formal congerence proofs [43, Section 5], which in turn represents an instance of the general program of formal proofs with explicit computational content proposed by Koutsoukou-Argyraki [35]. The authors plan to implement the results of this paper in the Lean [^2] https://www.lean-lang.org/ proof assistant, with the aim of not only formalising some central proof techniques in stochastic optimization, but providing programs that represent the computational content of those proofs, along with a range of techniques for generating such programs in a semi-automated way. This would both complement and build upon the growing literature on formalised stochastic algorithms, including the recent formalisations of Dvoretsky’s theorem [56] and the Doob martingale convergence theorems [60].
Acknowledgements. The authors are indebted to Nicholas Pischke for numerous insightful discussions and comments on the topics of this paper. The first author was partially supported by the EPSRC Centre for Doctoral Training in Digital Entertainment EP/L016540/1, and the second author was partially supported by the EPSRC grant EP/W035847/1.
Appendix A Proof of Theorem 4.2
Though already established in a much more general terms [44], in this appendix we prove Theorem 4.2 from first principles, taking advantage of the lack of generality to give a streamlined proof. We start with a variant of Doob’s well-known upcrossing inequalities [8] for martingales. Fixing for the remainder of the section a nonnegative supermartingale $\{{U_{n}}\}$, for $0\leq a<b$ let the random variable $C_{N}[a,b]$ denote the number of times $\{{U_{n}}\}$ crosses the interval $[a,b]$ up to time $N$, and similarly, $D_{N}[a,b]$ the number of times it downcrosses the interval. Since between any pair of downcrossings there must be exactly one upcrossing, we have $C_{N}[a,b]\leq 2D_{N}[a,b]+1$. Define $C[a,b]=\lim_{N\to\infty}C_{N}[a,b]$ and $D[a,b]$ similarly, noting that also $C[a,b]\leq 2D[a,b]+1$.
Lemma A.1.
Let $0\leq a<b$. Then
$\mathbb{E}[C[a,b]]\leq\frac{2\mathbb{E}[U_{0}]}{b-a}+1$
Proof.
The proof follows the standard pattern for deriving crossing inequalities (see e.g. [8] or [58]). Define the predictable process $Z_{n}\in\{0,1\}$ as follows: $Z_{1}=1$ if $U_{0}>b$ and $Z_{1}=0$ otherwise, and $Z_{n+1}=1$ if either $Z_{n}=1$ and $U_{n}\geq a$, or $Z_{n}=0$ and $U_{n}>b$, and if neither hold then $Z_{n+1}=0$. In other words, $\{{Z_{n}}\}$ represents a betting strategy for maximising our losses, where we start betting each time $\{{U_{n}}\}$ rises above $b$, and stop every time it falls below $a$. Our loss at time $N$ is then at least $(b-a)D_{N}[a,b]$ minus an error term for the current downcrossing not yet completed, or to be more precise:
(7) $\sum_{i=1}^{N}Z_{i}(U_{i}-U_{i-1})\leq-(b-a)D_{N}[a,b]+(U_{N}-b)^{+}$
Now let $Z^{\prime}_{n}:=1-Z_{n}$, which is also a predictable process, and since $\{{U_{n}}\}$ is a supermartingale we have
$\mathbb{E}\left[\sum_{i=1}^{N}Z^{\prime}_{i}(U_{i}-U_{i-1})\right]\leq 0$
for all $n\in\mathbb{N}$, and thus
$\mathbb{E}[U_{N}]-\mathbb{E}[U_{0}]$ $=\mathbb{E}\left[\sum_{i=1}^{N}(U_{i}-U_{i-1})\right]$
$\leq\mathbb{E}\left[\sum_{i=1}^{N}Z_{i}(U_{i}-U_{i-1})\right]+\mathbb{E}\left[\sum_{i=1}^{N}Z^{\prime}_{i}(U_{i}-U_{i-1})\right]$
$\leq\mathbb{E}\left[\sum_{i=1}^{N}Z_{i}(U_{i}-U_{i-1})\right]$
and so substituting the above into (7) we get
$(b-a)\mathbb{E}\left[D_{N}[a,b]\right]\leq\mathbb{E}[U_{0}]-\mathbb{E}[U_{N}]+\mathbb{E}[(U_{N}-b)^{+}]\leq\mathbb{E}[U_{0}]$
and thus
(8) $\mathbb{E}[D[a,b]]\leq\frac{\mathbb{E}[U_{0}]}{b-a}$
An alternative derivation of the same inequality for nonnegative supermartingales can be found in [9] (where (8) above follows immediately from (3.6) of that paper). The result now follows from $\mathbb{E}[C[a,b]]\leq 2\mathbb{E}[D[a,b]]+1$. ∎
Proof of Theorem 4.2.
Fix $\lambda,\varepsilon\in(0,1)$. First note that by Ville’s inequality (Lemma 4.1), for $K>\mathbb{E}[U_{0}]$ we have
$\mathbb{P}\left(|U_{n}|\geq\frac{2K}{\lambda}\right)\leq\mathbb{P}\left(\sup_{n\in\mathbb{N}}|U_{n}|\geq\frac{2K}{\lambda}\right)<\frac{\lambda}{2}$
and therefore it suffices to find a bound $\phi(\lambda,\varepsilon)$ such that for any $a_{0}<b_{0}\leq a_{1}<b_{1}\leq\ldots$ there exists some $n\leq\phi(\lambda,\varepsilon)$ satisfying
$\mathbb{P}\left(\exists k,l\in[a_{n};b_{n}]\left(|U_{k}-U_{l}|\geq\varepsilon\right)\cap|U_{a_{n}}|<\frac{2K}{\lambda}\right)<\frac{\lambda}{2}$
Fixing some arbitrary $a_{0}<b_{0}\leq a_{1}<b_{1}\leq\ldots$, define $Q_{n}$ to be the event inside the probability measure above. Divide $[0,2K/\lambda]$ into $p:=\lceil 8K/\lambda\varepsilon\rceil$ subintervals which we label $[\alpha_{j},\beta_{j}]$ for $j=0,\ldots,p-1$, and add a further interval of the same width at the upper end of $[0,2K/\lambda]$, which we label $[\alpha_{p},\beta_{p}]$. Each interval has width $42K/\lambda p\leq\varepsilon/$ 4.
Now take $\omega\in Q_{i}$, so that there exists $k(\omega),l(\omega)\in[a_{i};b_{i}]$ with $|U_{k(\omega)}(\omega)-U_{l(\omega)}(\omega)|\geq\varepsilon$ and also $|U_{a_{i}}(\omega)|<2K/\lambda$. By the triangle inequality, either $|U_{a_{i}}(\omega)-U_{k(\omega)}(\omega)|\geq\varepsilon/2$ or $|U_{a_{i}}(\omega)-U_{l(\omega)}(\omega)|\geq\varepsilon/2$. Since $U_{a_{i}}(\omega)\in[0,2K/\lambda]$ and we have our additional interval $[\alpha_{p},\beta_{p}]$, it follows that one of the intervals $[\alpha_{j},\beta_{j}]$ for $j=0,\ldots,p$ is crossed by $\{{U_{n}(\omega)}\}$ somewhere in $[a_{i};b_{i}]$, and therefore defining
$T_{i,j}:=\mbox{$\{{U_{n}}\}$ crosses $[\alpha_{j},\beta_{j}]$ somewhere in $[a_{i};b_{i}]$}$
we have shown that
$Q_{i}\subseteq\bigcup_{j=0}^{p}T_{i,j}$
Let $s_{N}:=\sum_{i=0}^{N}\mathbb{P}(Q_{i})$. Then
$s_{N}\leq\sum_{i=0}^{N}\mathbb{P}\left(\bigcup_{j=0}^{p}T_{i,j}\right)\leq\sum_{i=0}^{N}\sum_{j=0}^{p}\mathbb{P}\left(T_{i,j}\right)$
and so there is some $j\in\{0,\ldots,p\}$ such that
$\frac{s_{N}}{p+1}\leq\sum_{i=0}^{N}\mathbb{P}\left(T_{i,j}\right)=\mathbb{E}\left[\sum_{i=0}^{N}I_{T_{i,j}}\right]\leq\mathbb{E}\left[C[\alpha_{j},\beta_{j}]\right]$
Since $\beta_{j}-\alpha_{j}=2K/p\lambda$, by Lemma A.1 and using $\mathbb{E}[U_{0}]<K$ we have
$\mathbb{E}\left[C[\alpha_{j},\beta_{j}]\right]\leq\frac{2p\lambda\mathbb{E}[U_{0}]}{2K}+1=p\lambda+1$
Putting together the previous two inequalities and using $p<9K/\lambda\varepsilon$ we have
$s_{N}\leq(p+1)\left(p\lambda+1\right)<\left(\frac{9K}{\lambda\varepsilon}+1\right)\left(\frac{9K}{\varepsilon}+1\right)<\frac{100K^{2}}{\lambda\varepsilon^{2}}$
Finally, then, if $\mathbb{P}(Q_{i})\geq\lambda/2$ for all $i\leq N$ then
$\frac{(N+1)\lambda}{2}\leq s_{N}<\frac{100K^{2}}{\lambda\varepsilon^{2}}$
a contradiction for $N:=200K^{2}/\lambda^{2}\varepsilon^{2}$, which gives us our desired bound. ∎